Fairfax Pizza is considering buying a new oven. The new oven would be purchased today for 17,000 dollars. It would be depreciated straight-line to 1,600 dollars over 2 years. In 2 years, the oven would be sold for an after-tax cash flow of 2,100 dollars. Without the new oven, costs are expected to be 11,500 dollars in 1 year and 17,500 in 2 years. With the new oven, costs are expected to be 300 dollars in 1 year and 14,600 in 2 years. If the tax rate is 50 percent and the cost of capital is 11.34 percent, what is the net present value of the new oven project?
Answer #1

rate positively ..

Annual depreciation = (17000-1600)/2 7700
Computation of annual saving 1 2
With oven cost 300 14600
With out over cost 11500 17500
Saving = 11200 2900
year 0 1 2
A Initial investment -19600
operating cash flow
i Saving in cost      11,200           2,900
ii depreciation        7,700           7,700
iii=i-ii Taxable income        3,500         (4,800)
iv=iii*50% Tax@ 50%        1,750         (2,400)
v=iii-iv Profit after tax        1,750         (2,400)
B=v+ii operating cash flow after tax        9,450           5,300
C Post tax salvage value =           2,100
D=A+B+C Net cash flow after tax        (19,600)        9,450           7,400
E= PIVF @ 11.34% 1 0.89815 0.8066731
F=D*E Present value = (19,600.00) 8,487.52     5,969.38 (5,143.10)
Therefore NPV =     (5,143.10)